OpenStudy (yaboy):

Part A: Maria rented a coat at $285 for 3 days. If she rents the same coat for 6 days, she has to pay a total rent of $510. Write an equation in the standard form to represent the total rent (y) that Maria has to pay for renting the coat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (yaboy):

yes

OpenStudy (yaboy):

well if you can solve it for me it would be nice cause i have no clue

OpenStudy (3mar):

@Yaboy With @skullpatrol, you are in good hands!

OpenStudy (3mar):

No problem. That is a fact and I mean it!

OpenStudy (sshayer):

she pays some security and then rent for each day. in 3 days she pays security and rent for 3 days. In 6 days she pays same security and rent for 6 days. she pays rent for 3 days=510-285=225 security=285-225=60 so rent for one day=225/3=75 now please show me the total amount she pays for x days

OpenStudy (sshayer):

if she has no clue then we should give her some clue otherwise she will leave it or she may leave openstudy .